# Warm-Starting and Quantum Computing: A Systematic Mapping Study

### v2

@@ 3.3. Data Extraction and Analysis
RQ1 (publication trends) • Publication date e.g., June 2020
@@
• Publication venue e.g., IOP Quantum Science and Technology (Journal)
@@
• Type of research $\in$ {Validation Research, Evaluation Research, Solution Proposal, Philosophical Paper, Opinion Paper, Experience Paper}\} (Petersen et al., 2008; Wieringa et al., 2006)
RQ2 (algorithms & benefits) • Quantum algorithm(s) e.g., QAOA, VQE, QNN
• Computational problem(s) e.g., Maximum Cut, Portfolio Optimization, Classification
• Stated goals or benefits of WS e.g., speedup, accuracy increase, reduction of resource consumption
RQ3 (techniques) • Keywords e.g., quantum circuit, annealer, parameter initialization, initial state, classical-to-quantum, machine learning
@@ 4.3.1. Computational Problems and Quantum Algorithms in Warm-Starts
Classification 33 P09-T1; P09-T2; P09-T3; P10; P11; P12; P18; P23; P28; P29; P30; P31; P32; P33; P35; P36; P37; P38; P40; P42; P45; P51; P53; P56; P57; P59; P60; P62; P65; P66; P70; P71; P78
Maximum Cut 26 P01; P03; P04; P05; P06; P13; P14; P15; P16; P17; P21; P25; P27; P34; P41; P44; P46; P47; P48; P49; P55-T1; P55-T2; P72-T1; P72-T2; P77; P79
Image Classification 25 P09-T1; P10; P11; P18; P28; P30; P31; P33; P35; P37; P38; P40; P42; P45; P51; P53; P56; P59; P60; P62; P65; P66; P70; P71; P78
Ground State Estimation 8 P03; P08-T1; P08-T2; P14; P24; P26; P27; P80
TFIM (Traverse Field Ising Model) 4 P58-T1; P58-T2; P76-T1; P76-T2
Ground State Preparation 3 P20; P58-T1; P58-T2
Molecular Hamiltonian 2 P50; P61
Quantum State Classification 2 P09-T2; P09-T3
QUBO (Quadratic Unconstrained Binary Optimization) 2 P52; P63
Sherrington-Kirkpatrick Ising Models 2 P17; P34
3-SAT 1 P19
Eigenvalue 1 P22
Encoding 1 P02
Graph Bisection 1 P80
Knapsack Problem 1 P74
MAX-2-SAT 1 P80
MAX-3-SAT 1 P75
Modularity Maximization 1 P69
Molecular Geometry Optimization 1 P39
Portfolio Optimization 1 P04
Prime Factorization 1 P07
Transmon Simulation 1 P50
Variational State Preparation 1 P43
XXZ Spin Chain 1 P50
@@
QAOA Quantum Approximate Optimization Algorithm 35 P01; P03; P04; P05; P06; P13; P14; P15; P16; P17; P20; P21; P25; P34; P41; P43; P44; P46; P47; P48; P49; P55-T1; P55-T2; P63; P64; P69; P72-T1; P73-T1; P73-T2; P74; P76-T1; P76-T2; P77; P79; P80
QNN Quantum Neural Network 30 P02; P09-T1; P09-T2; P09-T3; P10; P11; P12; P18; P23; P28; P30; P31; P36; P37; P38; P39; P40; P42; P45; P51; P53; P56; P57; P59; P60; P62; P65; P70; P71; P78
VQE Variational Quantum Eigensolver 11 P03; P17; P23; P24; P26; P27; P50; P54; P61; P68; P80
HVA Hamiltonian Variational Ansatz 4 P08-T1; P08-T2; P58-T1; P58-T2
VQC Variational Quantum Circuit 4 P33; P35; P38; P40
HEA Hardware-Efficient Ansatz 3 P08-T1; P08-T2; P39
VQA Variational Quantum Algorithm 3 P27; P29; P67
AQC Adiabatic Quantum Algorithm 2 P07; P19
PQC Parameterized Quantum Circuit 3 P03; P14; P32
QA Quantum Annealing 2 P52; P72-T2
TQA Trotterized Quantum Annealing 1 P13
QPE Quantum Phase Estimation 1 P22
RBM Restricted Boltzmann Machine 1 P66
QCBM Quantum Circuit Born Machine 1 P68
CTQW Continuous-Time Quantum Walk 1 P75
QBM Quantum Boltzmann Machine 1 P39
@@ 4.4.1. Properties of Warm-Starting Techniques
Direction • C2Q Classical-to-Quantum: A quantum algorithm is warm-started utilizing classically generated information. 49
• Q2Q Quantum-to-Quantum: A quantum algorithm is warm-started utilizing information generated by a quantum algorithm. 37
• Q2C Quantum-to-Classical: A classical algorithm is warm-started utilizing information generated by a quantum algorithm. 2
Quantum Device • Annealer A quantum annealer is used. 5
• Gate-Based A gate-based quantum computer is used. 82
• Both Both kinds of devices are used. 1
Entry Point • Encoding in initial quantum state Information is encoded into the initial quantum state to conduct a warm-start. 37
• Quantum circuit construction The warm-start affects the circuit construction. 2
• Variational parameter initialization Generating initial parameters for VQAs. 36
• Annealing schedule Generating an annealing schedule. 4
• Optimization procedure Enhancing the optimization procedure. 7
• Encoding in classical neural network Encoding information into a classical neural network. 1
• Parameter initialization in classical neural network Generating parameters to initialize a classical neural network. 1
Change of Quantum Algorithm • Yes A quantum algorithm is adapted for the warm-start, i.e., a quantum circuit in the gate-based scenario. 14
• No No changes of quantum algorithms. 74
Reflexivity • Reflexive Information utilized to warm-start the targeted algorithm is generated by an instance of the same algorithm. 34
• Non-Reflexive Information is generated differently. 54
Data Generation • Only targeted problem instance required Only the targeted problem instance is needed to conduct the warm-start, including in any pre-processing. 35
• Data from other problem instances required Additional problem instances are solved to enable a warm-start. 53
@@
• Machine Learning (ML) Information is processed using ML to conduct a warm-start, e.g., to learn a model. 40
• Quantum Machine Learning (QML) Information processing by QML. 5
• None No or only trivial (classical) processing. 43

### v3

@@ 3.3. Data Extraction and Analysis
RQ1 (publication trends)
• Publication date
e.g., June 2020
• Type of publication venue
@@
• Publication venue
e.g., IOP Quantum Science and Technology (Journal)
• Author affiliations
@@
• Type of research
[…]
RQ2 (algorithms & benefits)
• Quantum algorithm(s)
e.g., QAOA, VQE, QNN
• Computational problem(s)
e.g., Maximum Cut, Portfolio Optimization, Classification
• Stated goals or benefits of WS
e.g., speedup, accuracy increase, reduction of resource consumption
RQ3 (techniques)
• Keywords
e.g., quantum circuit, annealer, parameter initialization, initial state, classical-to-quantum, machine learning
@@ 4.3.1. Computational Problems and Quantum Algorithms in Warm-Starts
Classification
33 P09-T1; P09-T2; P09-T3; P10; P11; P12; P18; P23; P28; P29; P30; P31; P32; P33; P35; P36; P37; P38; P40; P42; P45; P51; P53; P56; P57; P59; P60; P62; P65; P66; P70; P71; P78
Maximum Cut
26 P01; P03; P04; P05; P06; P13; P14; P15; P16; P17; P21; P25; P27; P34; P41; P44; P46; P47; P48; P49; P55-T1; P55-T2; P72-T1; P72-T2; P77; P79
Image Classification
25 P09-T1; P10; P11; P18; P28; P30; P31; P33; P35; P37; P38; P40; P42; P45; P51; P53; P56; P59; P60; P62; P65; P66; P70; P71; P78
Ground State Estimation
8 P03; P08-T1; P08-T2; P14; P24; P26; P27; P80
TFIM (Traverse Field Ising Model)
4 P58-T1; P58-T2; P76-T1; P76-T2
Ground State Preparation
3 P20; P58-T1; P58-T2
Molecular Hamiltonian
2 P50; P61
Quantum State Classification
2 P09-T2; P09-T3
QUBO (Quadratic Unconstrained Binary Optimization)
2 P52; P63
Sherrington-Kirkpatrick Ising Models
2 P17; P34
3-SAT
1 P19
Eigenvalue
1 P22
Encoding
1 P02
Graph Bisection
1 P80
Knapsack Problem
1 P74
MAX-2-SAT
1 P80
MAX-3-SAT
1 P75
Modularity Maximization
1 P69
Molecular Geometry Optimization
1 P39
Portfolio Optimization
1 P04
Prime Factorization
1 P07
Transmon Simulation
1 P50
Variational State Preparation
1 P43
XXZ Spin Chain
1 P50
@@
QAOA
Quantum Approximate Optimization Algorithm
35 P01; P03; P04; P05; P06; P13; P14; P15; P16; P17; P20; P21; P25; P34; P41; P43; P44; P46; P47; P48; P49; P55-T1; P55-T2; P63; P64; P69; P72-T1; P73-T1; P73-T2; P74; P76-T1; P76-T2; P77; P79; P80
QNN
Quantum Neural Network
30 P02; P09-T1; P09-T2; P09-T3; P10; P11; P12; P18; P23; P28; P30; P31; P36; P37; P38; P39; P40; P42; P45; P51; P53; P56; P57; P59; P60; P62; P65; P70; P71; P78
VQE
Variational Quantum Eigensolver
11 P03; P17; P23; P24; P26; P27; P50; P54; P61; P68; P80
HVA
Hamiltonian Variational Ansatz
4 P08-T1; P08-T2; P58-T1; P58-T2
VQC
Variational Quantum Circuit
4 P33; P35; P38; P40
HEA
Hardware-Efficient Ansatz
3 P08-T1; P08-T2; P39
VQA
Variational Quantum Algorithm
3 P27; P29; P67
AQC
Adiabatic Quantum Algorithm
2 P07; P19
PQC
Parameterized Quantum Circuit
3 P03; P14; P32
QA
Quantum Annealing
2 P52; P72-T2
TQA
Trotterized Quantum Annealing
1 P13
QPE
Quantum Phase Estimation
1 P22
RBM
Restricted Boltzmann Machine
1 P66
QCBM
Quantum Circuit Born Machine
1 P68
CTQW
Continuous-Time Quantum Walk
1 P75
QBM
Quantum Boltzmann Machine
1 P39
@@ 4.4.1. Properties of Warm-Starting Techniques
Direction
• C2Q
Classical-to-Quantum: A quantum algorithm is warm-started utilizing classically generated information. 49
• Q2Q
Quantum-to-Quantum: A quantum algorithm is warm-started utilizing information generated by a quantum algorithm. 37
• Q2C
Quantum-to-Classical: A classical algorithm is warm-started utilizing information generated by a quantum algorithm. 2
Quantum Device
• Annealer
A quantum annealer is used. 5
• Gate-Based
A gate-based quantum computer is used. 82
• Both
Both kinds of devices are used. 1
Entry Point
• Encoding in initial quantum state
Information is encoded into the initial quantum state to conduct a warm-start. 37
• Quantum circuit construction
The warm-start affects the circuit construction. 2
• Variational parameter initialization
Generating initial parameters for VQAs. 36
• Annealing schedule
Generating an annealing schedule. 4
• Optimization procedure
Enhancing the optimization procedure. 7
• Encoding in classical neural network
Encoding information into a classical neural network. 1
• Parameter initialization in classical neural network
Generating parameters to initialize a classical neural network. 1
Change of Quantum Algorithm
• Yes
A quantum algorithm is adapted for the warm-start, i.e., a quantum circuit in the gate-based scenario. 14
• No
No changes of quantum algorithms. 74
Reflexivity
• Reflexive
Information utilized to warm-start the targeted algorithm is generated by an instance of the same algorithm. 34
• Non-Reflexive
Information is generated differently. 54
Data Generation
• Only targeted problem instance required
Only the targeted problem instance is needed to conduct the warm-start, including in any pre-processing. 35
• Data from other problem instances required
Additional problem instances are solved to enable a warm-start. 53
@@
• Machine Learning (ML)
Information is processed using ML to conduct a warm-start, e.g., to learn a model. 40
• Quantum Machine Learning (QML)
Information processing by QML. 5
• None
No or only trivial (classical) processing. 43